OpenStudy (anonymous):

Gain a medal! Bill wants to factor x2 + 8x + 16 by grouping; however, Sara says it is a special product and can factor a different way. Using complete sentences, explain and demonstrate how both methods will result in the same factors.

OpenStudy (anonymous):

first off, lets factor, x2 + 8x + 16, by grouping ourselves x2+8x+16 x2+ 4x + 4x + 16 (x2+ 4x) + (4x + 16) x(x+4) + 4(x + 4) (x+4)(x+4) (x+4)^2 now lets factor by a special product x2+8x+16 a^2 + 2ab + b^2 = (a+b)^2 so, a=1 b=4 do you see how I picked out 'a' and 'b'?